Suppose that the economy of Stormwind has the following​ features:
The level of private saving is ​$150.
The level of investment expenditure is ​$150.
The level of exports is ​$100.
The level of imports is ​$150.
If​ Stormwind's economy is in​ equilibrium, the level of public saving would be equal to ​$
nothing. ​(Round your response to the nearest​ dollar.)
The level of public saving shows that the government in Stormwind has a budget

deficit
surplus
.
If the level of government purchases is ​$150 in​ Stormwind, the level of taxes must be ​$
nothing. ​(Round your response to the nearest​ dollar.)
1
Expert's answer
2019-10-03T09:12:30-0400

Private saving Spr = Y - T - C, public saving is Sp = T - G.

In equilibrium I = S, where S = Spr + Sp, so:

150 = 150 + Sp, so Sp = 0.

The level of public saving shows that the government in Stormwind has neither budget surplus, nor deficit.

If the level of government purchases is ​$150 in​ Stormwind, the level of taxes must be ​150$ too.